Presentation is loading. Please wait.

Presentation is loading. Please wait.

EMC Status Report: November 2003 Roger Barlow Events since last report (25 th Sept.) –The ROD Current status –Electronics –Neutron Generator –Light Pulser.

Similar presentations


Presentation on theme: "EMC Status Report: November 2003 Roger Barlow Events since last report (25 th Sept.) –The ROD Current status –Electronics –Neutron Generator –Light Pulser."— Presentation transcript:

1 EMC Status Report: November 2003 Roger Barlow Events since last report (25 th Sept.) –The ROD Current status –Electronics –Neutron Generator –Light Pulser –Chillers –Data Quality -Trickle Injection Christmas plans Running in 2004

2 ADB happenings 4 dead ADBs (Sept 25) became 6 (+1 noisy CARE) 3 in FB, 1 in BB, 2 in EC The ROD –Forward Barrel : Fixed 3 ADBs. –Endcap: correct pipe in East, assess in West –Wahid moved way up the learning curve –Attempt to fix Endcap East by IOB replacement. Fail (and lose neighbour) Score -1 –backward barrel left alone (door). Spontaneous cure. Score +1 Replaced FB ADB became noisy Now have 4 dead/noisy ADBs

3 Readout status Data from latest Light Pulser calibration run

4 Current Status: Electronics Dans next trip being organised for Jan/Feb 2004 FavouriteOnly OK ADB66 Barrel IOB11 Endcap IOB1?2 LV power supplies 5- Crates0- Crate power supplies 5-

5 ROB: Install Neutron Generator Its here. It works. Its paid for. Source calibration takes 30 Go back to 10 5 day between calibrations –Use `naturally occurring down times –Detailed instructions for shift crew to page Source Shifter Small work needed –Flashing blue light for operation during access –LV power supply surge protect and remote on/off control

6 Current Status: Light Pulser Used without airconditioning in cabinets (instability) Used for detecting dead readout in global detector calibration and independently by EMC New units on order, here December for installation over shutdown

7 Current Status: Worries(1) Endcap chiller. Not working (pump was dead, heating element for bypass line has problems + leaky valve…) and not understood. Leaking freon. Backup being used.

8 Re-engineering fan units in LV PS crates is going forward slowly/not at all. Currently 3 out of 12 crates done. There are no spare crates. If When one dies a replacement will have to be put together on the spot from spare parts. Current Status: Worries(2)

9 Current Status: Data Quality Run 4 looks OK. Trickle injection indications are that low background data with reasonable inhibit window are fine

10 Mass plots Run – normalRun – trickle

11 Christmas plans Replace noisy barrel ADB with a favourite Replace/reconnect Endcap East IOB Look at Backward Barrel miracle cure, replace a couple of ADBs with intermittent noisy CARE chips. Dont go for Endcap West. Pipe problems. Will need time and thought and expertise. Not worth risk of breaking something… Take time. Use trusted components.

12 Outlook for 2004 Anticipate starting with 1 dead ADB, to be fixed when we have plenty of time to work. Anticipate/hope for reasonably low failure rate so need for RODs is reasonably low: 3-4/year.


Download ppt "EMC Status Report: November 2003 Roger Barlow Events since last report (25 th Sept.) –The ROD Current status –Electronics –Neutron Generator –Light Pulser."

Similar presentations


Ads by Google